Hello, Tutors!
My name is D'Angelo. I was wondering can someone help me understand graphs/equations?
In my math class we are on this section in the book where its talking about graphs, and how to do equations. But, I don't understand because my teacher talks to fast, and you can bearly understand what she is saying so someone please help me.
I brought an example of what where doing, but I don't know how to do them.
y=-2x + 6
y=x - 4
Thank You
D'Angelo
Thank You

Answer by sabanasir(22)

well its not a problem.... d'nt panic :)
u take a range ok...
lets take
-3,-2,-1,0,1,2,3
no remember, that it shud b always a series of real numbers. if u want to start from -4 then it should extend to 4... Get it?
now what we shall do is:
take the above range and start putting them into equations of y= -2x+6 and y= x-4
where we'll put these values??? we shall put them one by one at 'x'.
like:
consider the 1st equation:
y = -2x + 6
when x=-3
y = -2(-3) + 6
=> y = 6 + 6=12
no the first coordinate for this equation is= (x,y)=(-3,12) meaning you will plot a point where x is -3 and along with it y is 12... get it??
yopu have to count the squares on either side.. on -X side count till 3 then see where 12 comes on Y line. Get it??
simililary solve for y where 'x' is -2,-1,0,1,2,3 then when u get all your point for these values of 'x' you graph them up and join them.
u shall be following the exact same method for the second equation as well.Good Luck